The following set of lead determinations on separate aliquots of a pooled blood serum were reported: 103,106,107 and 114 meq/ L. One of the values appears suspect. Should the suspected be retained or discarded at 95% confidence level?

1
Expert's answer
2022-05-20T12:51:02-0400

Let use Q test.

Q=gap/range

gap=114-107=7

range=114-103=11

Q=7/11=0.636

Qcritic=0.829

0.636<0.829

Q<Qcritic

The suspected be retained at 95% confidence level.